OpenStudy (anonymous):

Rationalize the denominator: √50 over √5

OpenStudy (unklerhaukus):

\[\frac{\sqrt{50}}{\sqrt 5}\]

OpenStudy (unklerhaukus):

\[\sqrt{ab}=\sqrt a\times\sqrt b\]

OpenStudy (anonymous):